Two factory plants are making TV panels. Yesterday, Plant A produced 7000 fewer panels than plant B did. Five percent of the panels from plant A and 3% of the panels from plant B were defective. How many of panels did plant B produce, if the two plants together produced 1010 defective panels.
Answer by ikleyn(52846)

.
Two factory plants are making TV panels.
Yesterday, Plant A produced 7000 fewer panels than plant B did.
Five percent of the panels from plant A and 3% of the panels from plant B were defective.
How many of panels did plant B produce, if the two plants together produced 1010 defective panels.
~~~~~~~~~~~~~~~~~~~

Let x = # of TV panel produced by plant B.

Then the number of TV panels produced by plant A is (x-7000).


From the problem's description, you have this equation

    0.05(x-7000) + 0.03*x = 1010  defective panel.


Simplify and find x

    0.05x - 350 + 0.03x = 1010

            0.08x       = 1010 + 350

            0.08x       =    1360

                x       =    1360/0.08 = 17000.


ANSWER.  Plant B produced 17000 TV panels.

Solved.